OpenStudy (anonymous):

Which ratio is equivalent to 8 : 32? A. 1 : 16 B. 1 : 4 C. 1 : 24 D. 1 : 8

OpenStudy (anonymous):

@DanJS

pooja195 (pooja195):

Divide both numbers by 8 you need to find a common number that can divide 8 is perfect 8/8=? 32/8=?

OpenStudy (danjs):

ratio is a fraction too 8 /32 , then same as last prob

OpenStudy (anonymous):

I think it's B

pooja195 (pooja195):

Correct

563blackghost (563blackghost):

Your correct ^^

OpenStudy (anonymous):

8/8 is 1 and 32/8 is 4

pooja195 (pooja195):

correct you have the answer doe ;p

OpenStudy (anonymous):

Okay thanks guys

OpenStudy (danjs):

both numbers factor into a bunch of 2's \[\huge \frac{ 2^3 }{ 2^5 }\]

OpenStudy (anonymous):

thats really helpful @DanJS

OpenStudy (danjs):

leaves 1/2^2 or 1/4